OpenStudy (anonymous):

Write the expression as one involving only sin theita and cos theita: sin2theita+cos3theita

OpenStudy (priyar):

do u know \[\sin2\theta \] expansion?

OpenStudy (mathmale):

Your "sin2theita+cos3theita" is open to misinterpretation and needs clarification. I suspect you mean\[\sin 2\theta + \cos 3\theta\] ... but some might interpret your sin2theta to be \[\sin^2\theta\] Please clarify. The best way to avoid ambiguity here would be to use parentheses around your "2 theta" or (sin theta)^2, whichever the case may be.

OpenStudy (mathmale):

You might find it helpful to recognize that\[\cos 3 \theta = \cos (2 \theta+1\theta)\] and to expand that \[\cos 2\theta \] using the proper double angle identity.